### v1

@@ 4 Structured Semantic Augmentation (SSA)
When merging two vgAMRs, the main challenge is identifying which nodes correspond to the same concepts, such as entities, attributes, actions, and relations. We use three node properties to accomplish this: a) visual grounding information, b) semantic similarity of node labels, and c) node neighborhood semantic similarity. We derive two node-merging criteria from there: 1) visually grounded entity nodes are merged if they point to the same image-bounding boxes. When 1) does not hold, we check the second criterion: 2) for the remaining non-grounded nodes, including amr-specific, predicates, adjectives, and adverbs, we use a combination of node label semantic similarity (cosine similarity of the labels using their GloVe embeddings) and neighborhood similarity. Neighborhood similarity examines the similarity of parents for adjectives/adverbs nodes and children for predicate nodes, along with the similarity of connecting edge roles. When two nodes satisfy criterion 1) or 2), we merge them into a single node. Moreover, if they have different labels, we maintain both names by keeping a list of synonyms to increase representation diversity. In Appendix 0.B (Fig. 7) we have included the flow diagram depicting the process of merging two nodes corresponding to the same concepts.
@@
We use the SPRING AMR-to-Text model [9] to generate new event-focused captions from the sampled vgAMR subgraphs. Because both vgAMR merging and sampling steps introduce noise, the output captions are not always of good quality. We automatically filter low-quality captions by using a linguistic well-formedness measure, GRUEN [63], which is a reference-free metric based on BERT contextual embeddings.
In Sec. 0.D.7 we provide examples of original dataset captions and their SSA augmentations, along with their GRUEN score.
@@ 4.1 Mixing Strategies of Original and SSA Data
Due to space limitations, we present results from the Random Sampling Strategy for $p_{SSA}=0\%$ and $p_{SSA}=100\%$ in the main paper. Results from other scenarios can be found in Sec. 0.D.2.
@@ 5.1 Data
We use Flickr30k Entities (Flickr-Ent) [40] and MS-COCO Entities (COCO-Ent) [19] for training and evaluation. Flickr-Ent augments the original captions of Flickr30k [56] with manually-annotated region–phrase groundings. Flickr-Ent contains the original $31K$ images annotated with five captions each. COCO-Ent augments the original MS-COCO [29] ($120K$ images each annotated with around five captions) with semi-automatically collected grounding annotations; see [19] for details on the annotation process. For both datasets, we follow previous work and use the training and test splits by Karpathy et al. [26].
We apply our SSA algorithm on the aforementioned datasets to create their augmented variations, COCO-Ent-SSA and Flickr-Ent-SSA, containing about $800K$ and $250K$ training captions, respectively, of which $33\%$ and $37\%$, are generated by our SSA algorithm.
@@
5.2 Models and Evaluation Metrics
We report the performance of our models and baselines using a comprehensive set of metrics that evaluate different aspects of caption controllability, diversity, and quality. We also propose and report an overall performance metric that summarizes these different aspects in a meaningful way. In particular, for diversity, we measure n-gram diversity D-1, D-2 [4] and self-CIDEr (sC) [51] metrics. For content controllability, we use an extended version of the IoU [19] metric and the proposed Hallucinations (Hal) metric based on IoU, described in detail in Sec. 0.C.1. For length controllability, we measure mean absolute error (L) and length precision (LP) [20]. We assess the generated text quality using the GRUEN (G) [63] metric. We determine the overall performance using the harmonic mean of IoU, G, and sC metrics. A higher score indicates better performance. The harmonic mean ($H$) helps us determine the model with the best overall performance since it prioritizes models that perform well across all metrics while penalizing those with poor performance, even in one metric. Sec. 0.C.1 provides details for each evaluation metric. Further, in Sec. 0.D.5 we include comparisons of the CIC models on standard captioning metrics (like CIDEr [49] and Spice [2]). Note that standard captioning metrics are not sufficient for evaluating CIC, as they compare a generated (controlled) caption with a reference ground-truth caption, ignoring the desired effect of the control signal. We report them for completeness, but we believe that our well-formedness metric (GRUEN) is more suited for evaluating the quality of the controlled captions.
@@ 6.1 Overall Performance
We compare the length precision of our model with the baselines utilizing length control in Table 3. LaBERT uses only length-control signals without spatial control, while our model employs both spatial and length-control signals to generate focused captions. This makes the LaBERT task much easier since it only focuses on generating specific length descriptions of an image. On the other hand, our model focuses on generating captions that describe only a specific sub-region of the scene while maintaining a desired description length level. Although our task is more challenging than LaBERT, we achieve competitive length precision performance. Lastly, we want to emphasize that the improvement in length controllability (L) and length precision (LP) from CIC-BART to CIC-BART-SSA stems from the increased length diversity found in our SSA augmentations, which enriches the original COCO-Ent and Flickr-Ent datasets.
In Sec. 0.D.1, we provide an analysis of the caption length statistics in the original datasets and our SSA-derived captions.
@@ 6.2 Effect of SSA on Content Controllability
In Sec. 0.D.3, we include the content hallucinations (Hal) performance across different levels of focus of the control signals.
@@ 6.3 Qualitative Analysis
In Fig. 5, we present qualitative examples from the original test sets. examples from our SSA (only) test set control signals. In these examples, the control signals are extracted from the ground-truth captions. In the two figures, each highlighted word found in the generated controlled captions corresponds to the control entity of the same color. This shows the match between the captions produced and the control signal. We also strike through the parts where the model hallucinates or generates redundant references to the entities of interest. Our models have been observed to outperform the previous state-of-the-art models by substantially enhancing the quality of the generated controlled captions. This behavior was expected from our quantitative analysis, which showed that our models have significantly higher text quality (G). More importantly, our CIC-BART-SSA model is capable of generating captions that are faithful to the control signal and better understand the relationships that connect the entities of interest.
@@
In addition, more evidently, in figures (b) and (d), our models have better content controllability performance by correctly referring to all entities of interest in the control signal. Especially in the highly challenging, complex scene (d) in which many objects are present, it successfully describes all entities of interest in the generated controllable caption. Our CIC-BART-SSA model generates captions that are faithful to the control signal and better understand the relationships connecting the entities of interest. For example, in (a), it correctly identifies that the person is photographing his food rather than eating it and that the image is not black and white, or in (d) that the woman holds the purse and not the man in the background.
We include additional qualitative samples in Sec. 0.D.6.
@@ Appendix 0.A Overview
In our supplementary material, we provide more details for our SSA methodology in Appendix 0.B; in Appendix 0.C, we provide additional details for our experimental set-up; and finally, in Appendix 0.D we provide extended qualitative and quantitative results of our performed experiments and ablations. Specifically, we focus on:
@@
Figure 7: The flow diagram shows the process of merging nodes A and B from two vgAMRs. The merging process starts with handling the AMR-specific nodes of type ’AND’ and ’OR’, followed by the visually grounded vgAMR nodes. For the remaining non-grounded nodes (such as predicates, adjectives, and nouns), we check if their names are very similar to decide if they refer to the same concepts and can be merged. If the nodes do not fit the above categories, we check if they are predicate nodes. If they are, we examine their neighborhood to determine if they can be merged.
@@ Appendix 0.B Structured Semantic Augmentation (SSA)
In this section, we will provide additional information on the SSA augmentation strategy we introduced in our main paper. We summarize the steps in constructing the meta-vg Graph (as described in Step 1: Image-level AMR graph generation in our main paper) in Algorithm 1. We present the flow diagram that explains the process of determining if two nodes from different vgAMRs refer to the same concept and, therefore, should be merged in Fig. 7. Furthermore, we include a detailed example of our SSA methodology in Fig. 8.
As illustrated in Fig. 8, we begin with five original visually grounded captions from an existing captioning dataset, which we transform into their AMR graphs using pre-trained Text-to-AMR parsers. Then, we introduce a new type of AMR, the visually grounded AMRs (vgAMRs), which we create using the AMR nodes-captions word alignments and the caption words’ visual grounding. We represent the grounding information in our vgAMRs by using colored node labels.
To construct the hierarchical clusters, we use the UPGMA algorithm, which considers each individual vgAMR as a separate cluster at Level 0. Two clusters are merged at each level based on their distance, starting with the most similar graphs. To measure similarity, we use the SmatchScore between two vgAMR graphs. Since the Smatch score is a metric from 0 to 1, we use 1- Smatch score as the distance metric for the UPGMA algorithm.
@@
For this example, the AMR graphs of captions (2) and (3) are the most similar, so they are merged first to create their joint vgAMR graph at Level 1. Every graph from levels 1 to 4 results from the 6-17 step of our Algorithm 1 where we merge two graphs from lower layers according to the hierarchical clusters computed by UPGMA. The final layer (4) graph is our meta-vgAMR graph, which contains all information from the original vgAMRs and, thus, from the available original captions.
By applying our event-focused sampling approach, we can generate novel, focused, visually grounded descriptions from this new structure. Some examples can be seen at the bottom of Fig. 8, along with the resulting captions generated by pre-trained AMR-To-Text parsers.
@@ 0.C.1 Evaluation Metrics
To measure content controllability we use an extended version of the IoU metric of [19] that calculates the degree-of-match (faithfulness) between a control signal and the corresponding generated caption. For our control signal, we use the set of nouns $\mathcal{E}$ that represent the entities of interest, which are the names of the visual objects in the control. To extract nouns from the predicted sentences, we use the Stanford part-of-speech tagger [47, 46]. We then find the semantic intersection of the two sets using Hungarian Matching, as in [19]. Finally, we calculate the semantic intersection over union of the control nouns and the nouns extracted from the controllable caption, which gives us our content controllability IoU.
@@
For a fair comparison of the different CIC models, we measure diversity for the five generated captions for each test image (in COCO-Ent and Flickr-Ent), and report their average. Note that not all images in COCO-Ent and Flickr-Ent have five caption–control signal pairs, especially for COCO-Ent that is automatically annotated. We only considered the ones with five available pairs for diversity evaluation, including $985$ images for Flickr-Ent and $112$ images for COCO-Ent. For a given image and $M$ randomly generated control signals, we select all possible sets of $5$ captions ($M$ choose $5$) and measure the ratio of $n$-grams to total number of words for each set. We report the average of the best Div-$n$ scores for all images in the test set.
@@ 0.C.2 CIC-BART-SSA and Baselines Setup
In Fig. 9, we present an example with the control signals used by the baselines and our models for a specific instance where we need a focused caption on the boy (bbox 1) and the cake (bbox2, bbox3). The SCT model [19] uses bounding boxes of entities of interest and GLoVE embeddings of their Faster R-CNN labels as control signals. The VSR model [13] adds ground truth caption verbs and their PropBank grounded verb semantic roles to the SCT control signal. The ASG model [14] employs abstract scene graphs as control signals that provide information about how visual entities are related or connected and how many attributes they have.
Our models (CIC-BART and CIC-BART-SSA) use only the bounding boxes of interest and the desired caption length level as control signals. We have also explored the use of ground truth verbs in the control signal, like in VSR, in our CIC-BART +verb model. However, unlike VSR, we only use the ground truth verb name and not their PropBank grounded semantic roles.
@@ 0.D.6 Qualitative Results
In Fig. 14, we present qualitative examples from COCO-Ent and Flickr-Ent test sets. In these examples, the control signals are extracted from the ground-truth captions. Each colored oval under “Cntl” corresponds to a bounding box of the same color in the image. The collection of ovals identifies the entities of interest, that is, the control signal. (Note that we do not show the colored ovals for image (c), since both sets of control signals include all bounding boxes.) For example, in (a), the first control signal (at the top) focuses on the regions restaurant, man, and food, while the second control signal also includes the entity table. Each highlighted word in the generated controlled captions corresponds to the control entity of the same color, showing the match between the generated captions and the control signal.

### v2

@@ Abstract
footnotetext: *Work done during an internship at Samsung AI Centre - Torontofootnotetext: †Work done while at Samsung AI Centre - Toronto
@@
Keywords:
Controllable Image Captioning Vision Language Model Abstract Meaning Representation
@@ 4 Structured Semantic Augmentation (SSA)
When merging two vgAMRs, the main challenge is identifying which nodes correspond to the same concepts, such as entities, attributes, actions, and relations. We use three node properties to accomplish this: a) visual grounding information, b) semantic similarity of node labels, and c) node neighborhood semantic similarity. We derive two node-merging criteria from there: 1) visually grounded entity nodes are merged if they point to the same image-bounding boxes. When 1) does not hold, we check the second criterion: 2) for the remaining non-grounded nodes, including amr-specific, predicates, adjectives, and adverbs, we use a combination of node label semantic similarity (cosine similarity of the labels using their GloVe embeddings) and neighborhood similarity. Neighborhood similarity examines the similarity of parents for adjectives/adverbs nodes and children for predicate nodes, along with the similarity of connecting edge roles. When two nodes satisfy criterion 1) or 2), we merge them into a single node. Moreover, if they have different labels, we maintain both names by keeping a list of synonyms to increase representation diversity. In Appendix 0.B (Fig. 8), we have included the flow diagram depicting the process of merging two nodes corresponding to the same concepts.
@@
We use the SPRING AMR-to-Text model [9] to generate new event-focused captions from the sampled vgAMR subgraphs. Because both vgAMR merging and sampling steps introduce noise, the output captions are not always of good quality. We automatically filter low-quality captions by using a linguistic well-formedness measure, GRUEN [63], which is a reference-free metric based on BERT contextual embeddings. In Sec. 0.D.7 we provide examples of original dataset captions and their SSA augmentations, along with their GRUEN score.
@@ 4.1 Mixing Strategies of Original and SSA Data
We present results from the Random Sampling Strategy for $p_{SSA}=0\%$ and $p_{SSA}=100\%$ in the main paper. Results from other scenarios can be found in Sec. 0.D.2.
@@ 5.1 Data
We use Flickr30k Entities (Flickr-Ent) [40] and MS-COCO Entities (COCO-Ent) [19] for training and evaluation. Flickr-Ent augments the original captions of Flickr30k [56] with manually-annotated region–phrase groundings. Flickr-Ent contains the original $31K$ images annotated with five captions each. COCO-Ent augments the original MS-COCO [29] ($120K$ images each annotated with around five captions) with semi-automatically collected grounding annotations; see [19] for details on the annotation process. For both datasets, we follow previous work and use the training and test splits by Karpathy et al. [26]. We apply our SSA algorithm on the aforementioned datasets to create their augmented variations, COCO-Ent-SSA and Flickr-Ent-SSA, containing about $800K$ and $250K$ training captions, respectively, of which $33\%$ and $37\%$, are generated by our SSA algorithm.
@@
5.2 Models and Evaluation Metrics
We report the performance of our models and baselines using a comprehensive set of metrics that evaluate different aspects of caption controllability, diversity, and quality. We also propose and report an overall performance metric that summarizes these different aspects in a meaningful way. In particular, for diversity, we measure n-gram diversity D-1, D-2 [4] and self-CIDEr (sC) [51] metrics. For content controllability, we have developed an extended version of the IoU [19] and further analyzed its performance by introducing the Hallucinating Nouns (Hal) metric. Both metrics are thoroughly discussed in Sec. 0.C.1. For length controllability, we measure mean absolute error (L) and length precision (LP) [20]. We assess the generated text quality using the GRUEN (G) [63] metric. We determine the overall performance using the harmonic mean of IoU, G, and sC metrics. A higher score indicates better performance. The harmonic mean ($H$) helps us determine the model with the best overall performance since it prioritizes models that perform well across all metrics while penalizing those with poor performance, even in one metric. Sec. 0.C.1 provides details for each evaluation metric. Further, in Sec. 0.D.5 we include comparisons of the CIC models on standard captioning metrics (like CIDEr [49] and Spice [2]). Note that standard captioning metrics are not sufficient for evaluating CIC, as they compare a generated (controlled) caption with a reference ground-truth caption, ignoring the desired effect of the control signal. We report them for completeness, but we believe that our well-formedness metric (GRUEN) is more suited for evaluating the quality of the controlled captions.
@@ 6.1 Overall Performance
We compare the length precision of our model with the baselines utilizing length control in Table 3. LaBERT uses only length-control signals without spatial control, while our model employs both spatial and length-control signals to generate focused captions. This makes the LaBERT task much easier since it only focuses on generating specific length descriptions of an image. On the other hand, our model focuses on generating captions that describe only a specific sub-region of the scene while maintaining a desired description length level. Although our task is more challenging than LaBERT, we achieve competitive length precision performance. Lastly, we want to emphasize that the improvement in length controllability (L) and length precision (LP) from CIC-BART to CIC-BART-SSA stems from the increased length diversity found in our SSA augmentations, which enriches the original COCO-Ent and Flickr-Ent datasets. In Sec. 0.D.1, we provide an analysis of the caption length statistics in the original datasets and our SSA-derived captions.
@@ 6.3 Qualitative Analysis
In Fig. 5, we present qualitative examples from the original test sets, and in Fig. 6 examples from our SSA (only) test set control signals. In the two figures, each highlighted word found in the generated controlled captions corresponds to the control entity of the same color. This shows the match between the captions produced and the control signal. We also strike through the parts where the model hallucinates or generates redundant references to the entities of interest.
@@
Our models have been observed to outperform the previous state-of-the-art models by substantially enhancing the quality of the generated controlled captions. This behavior was expected from our quantitative analysis, which showed that our models have significantly higher text quality (G). More importantly, our CIC-BART-SSA model is capable of generating captions that are faithful to the control signal and better understand the relationships that connect the entities of interest. We include additional qualitative samples in Sec. 0.D.6.
@@ Appendix 0.A Overview
In the appendix, we provide more details for our SSA methodology in Appendix 0.B; in Appendix 0.C, we provide additional details for our experimental set-up; and finally, in Appendix 0.D we provide extended qualitative and quantitative results of our performed experiments and ablations. Specifically, we focus on:
@@
Figure 8: The flow diagram shows the process of merging nodes A and B from two vgAMRs. The merging process starts with handling the AMR-specific nodes of type ‘AND’ and ‘OR’, followed by the visually grounded vgAMR nodes. For the remaining non-grounded nodes (such as predicates, adjectives, and nouns), we check if their names are very similar to decide if they refer to the same concepts and can be merged. If the nodes do not fit the above categories, we check if they are predicate nodes. If they are, we examine their neighborhood to determine if they can be merged.
@@ Appendix 0.B Structured Semantic Augmentation (SSA)
In this section, we will provide additional information on the SSA augmentation strategy we introduced in our main paper. We summarize the steps in constructing the meta-vg Graph (as described in Step 1: Image-level AMR graph generation in our main paper) in Algorithm 1. We include a detailed example of our SSA methodology in Fig. 7. We present the flow diagram that explains the process of determining if two nodes from different vgAMRs refer to the same concept and, therefore, should be merged in Fig. 8.
SSA Algorithm.
@@
To construct the hierarchical clusters, we use the UPGMA algorithm, which considers each individual vgAMR as a separate cluster at Level 0. Two clusters are merged at each level based on their distance, starting with the most similar graphs. To measure similarity, we use the Smatch Score between two vgAMR graphs. Since the Smatch score is a metric from 0 to 1, we use 1- Smatch score as the distance metric for the UPGMA algorithm. For this example, the AMR graphs of captions (2) and (3) are the most similar, so they are merged first to create their joint vgAMR graph at Level 1. Every graph from levels 1 to 4 results from the 6-17 step of our Algorithm 1 where we merge two graphs from lower layers according to the hierarchical clusters computed by UPGMA. The final layer (4) graph is our meta-vgAMR graph, which contains all information from the original vgAMRs and, thus, from the available original captions. By applying our event-focused sampling approach, we can generate novel, focused, visually grounded descriptions from this new structure. Some examples can be seen at the bottom of Fig. 7, along with the resulting captions generated by pre-trained AMR-To-Text parsers.
Finding same-concept nodes in two vgAMR graphs.
In the flow diagram labeled as Fig. 8, we can observe that when we merge two vgAMR graphs, vgAMR-A and vgAMR-B, we need to identify the common concept nodes between the two representations and combine them. This merging process serves two purposes: a) it allows for a more efficient and compressed representation by reducing redundancies and eliminating multiple nodes for the same concept, and b) it consolidates all available information about a particular concept found in different captions. For example, in Fig. 14 c) for the top player, one caption may describe her clothing, another her physical characteristics, and yet another her actions (for instance, practicing martial arts). Despite the differences, all these captions have a common concept: the person in the picture. Instead of having three separate nodes with partial information, we aim to create a single node (person) that consolidates all available information about the person in the image, making it easier to explore all connected nodes and access the complete information.
The process for identifying common nodes involves the steps outlined in Fig. 8. We start by checking if the two nodes are AMR-specific nodes of type AND. If these nodes are found at the root of the graph, it indicates that the corresponding sentence follows the format ‘FACT-1 AND FACT-2’. In this case, we can merge them as they represent aggregated facts about the image. If they are not root nodes, we need to be more cautious and ensure that they originate from the same concept. For this reason, we assess the similarity of their neighboring nodes. If they link to the same nodes, we merge them and combine the provided facts.
As shown in Fig. 8, if both nodes are visually grounded and refer to the same visual entity (i.e. if they have the same bounding boxes), we are hesitant to merge the nodes without first verifying that their names are synonyms  [^4] We determine if two nodes are synonyms by comparing the cosine similarity of their GloVe embeddings. If the cosine similarity is above a certain threshold, we consider them synonyms.. This additional condition is helpful in cases where a) the original dataset visually grounds phrases instead of nouns, and b) there is noise from the Text-to-AMR parser. This check ensures that entity attributes (such as ‘young’ and ‘tall’) which may be visually grounded, are not mistakenly merged with noun nodes. Finally, if the two nodes are semantically distant or do not refer to the same visual entity, or if one of them is grounded and the other is not, we conclude that the two nodes cannot be merged.
When we don’t have visual cues to help us identify similar concept nodes, we rely on the names of the nodes and their surrounding information to make decisions. If two nodes are synonyms, we look at how similar their neighbors are (e.g., if the two nodes are nouns or adjectives, do they share the same parent?). If they do, we merge them as similar concept nodes.
In our final step, we have an additional procedure for predicate nodes. In our experiments, we observed that the GloVe embeddings of predicate/verb words tend to be more distant. Therefore, in the last step, if the two nodes are predicates, and their child nodes (ARG0, ARG1, and so on) are the same, and the similarity of their names is above a certain threshold (which is smaller than the thresholds used in the previous steps), then we merge the two predicate nodes. This concludes our node merge process.
@@ 0.C.1 Evaluation Metrics
To measure content controllability we design an extended version of the IoU metric of [19] that calculates the degree-of-match (faithfulness) between a control signal and the corresponding generated caption. For our control signal, we use the set of nouns $\mathcal{E}$ that represent the entities of interest, which are the names of the visual objects in the control. To extract nouns from the predicted sentences, we use the Stanford part-of-speech tagger [47, 46]. We then find the semantic intersection of the two sets using Hungarian Matching, as in [19]. Finally, we calculate the semantic intersection over union of the control nouns and the nouns extracted from the controllable caption, which gives us our content controllability IoU.
@@ 0.C.2 CIC-BART-SSA and Baselines Setup
In Fig. 9, we present an example with the control signals used by the baselines and our models for a specific instance where we need a focused caption on the boy (bbox 1) and the cake (bbox2, bbox3). The SCT model [19] uses bounding boxes of entities of interest and GLoVE embeddings of their Faster R-CNN labels as control signals. The VSR model [13] adds ground truth caption verbs and their PropBank grounded verb semantic roles to the SCT control signal. The ASG model [14] employs abstract scene graphs as control signals that provide information about how visual entities are related or connected and how many attributes they have. Our models (CIC-BART and CIC-BART-SSA) use only the bounding boxes of interest and the desired caption length level as control signals. We have also explored the use of ground truth verbs in the control signal, like in VSR, in our CIC-BART +verb model. However, unlike VSR, we only use the ground truth verb name and not their PropBank grounded semantic roles.
@@ 0.D.6 Qualitative Results
In Fig. 14, we present qualitative examples from COCO-Ent and Flickr-Ent test sets. In these examples, the control signals are extracted from the ground-truth captions. Each colored oval under ‘Cntl’ corresponds to a bounding box of the same color in the image. The collection of ovals identifies the entities of interest, that is, the control signal. (Note that we do not show the colored ovals for image (c), since both sets of control signals include all bounding boxes.) For example, in (a), the first control signal (at the top) focuses on the regions restaurant, man, and food, while the second control signal also includes the entity table. Each highlighted word in the generated controlled captions corresponds to the control entity of the same color, showing the match between the generated captions and the control signal.